OpenStudy (anonymous):

what number lies between 4 and 4.5 on a number line? The choices are sqrt13, sqrt17, sqrt22, and sqrt28.

hartnn (hartnn):

means you need to find the mid-point between 4 an 4.5 just take the average! \(\Large \dfrac{4+4.5}{2} =... ?\)

OpenStudy (anonymous):

I didn't get any of the choices when I did that :/ have any more ideas to help me?

hartnn (hartnn):

strange choices, do you have any other similar solved example ?

OpenStudy (anonymous):

no, I don't have any that look like this

hartnn (hartnn):

ok, i think they need any number between 4 and 4.5 so we'll take each option one by one

hartnn (hartnn):

sqrt 13 where on number line you think this will lie ?

OpenStudy (anonymous):

if it helps this question is about real numbers and comparing them

hartnn (hartnn):

let me give you an example sqrt 28 since 5^2 is 25 sqrt 25 = 5 and sqrt 28 is just lil bit greater than sqrt 25 so, sqrt 28 will lie between 5 and 5.5 did you get this ?

OpenStudy (anonymous):